1. given the equation 5x-2y=6, find two points on the line.
( , )( , )

Answer by mananth(16949)

given the equation 5x-2y=6, find two points on the line.
( , )( , )
put x= 0 y=-3 (0,-3)
...
Put y=0 , x = 5/6 (5/6,0)